Kay rode the bicylce into the country at 10mph and Yancy pushed it back to town at 3 mph. If the round trip took 13 hours, how far did Kay ride the bicylce into the country?
Answer by checkley71(8403)

x/10+x/3=13
(3x+10x)/30=13
13x/30=13
13x=13*30
13x=390
x=390/13
x=30 miles is the mileage into the country.
proof.
30/10+30/3=13
3+10=13
13=13